The plane 3x+4y+2z=1 touches the conicoid 3x^2+2y^2+z^2=1
True or false with full explanation
1
Expert's answer
2020-12-17T06:41:19-0500

The plane "3x+4y+2z=1" touches the conicoid "3x^2+2y^2+z^2=1."

True or false

Solution:

Suppose that the plane

"3x+4y+2z=1"

touches the conicoid

"3x^2+2y^2+z^2=1"

at some point

"M(x_0,y_0,z_0)."

Then the plane equation can be written as follows:

"F_{x'}\mid_M(x-x_0)+F_{y'}\mid_M(y-y_0)+F_{z'}\mid_M(z-z_0)=0."

where

"F(x,y,z)=3x^2+2y^2+z^2-1,"

"F_{x'}\mid_M=6x_0,\space F_{y'}|_M=4y_0,\space F_{z'}|_M=2z_0"

or

"6x_0(x-x_0)+4y_0(y-y_0)+2z_0(z-z_0=0),"

"6x_0x+4y_0y+2z_0z-(6x_0^2+4y_0^2+2z_0^2)=0."

The coefficients of the last equation must be equal to the coefficients of the equation

"3x+4y+2z-1=0."

Check it out:

"6x_0=3,\space 4y_0=4,\space 2z_0=2,\space 6x_0^2+4y_0^2+2z_0^2=1."

"x_0=\frac{1}{2},\space y_0=1,\space z_0=1,\space 7.5=1."

The last equality is wrong,

Answer:

so the plane

"3x+4y+2z=1"

does not touch the conicoid

"3x^2+2y^2+z^2=1"

and there is no point of contact.